OpenStudy (anonymous):

Help please! Write x^2 - 8x + 13 = 0 in the form (x - a)^2 = b, where a and b are integers. (x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (anonymous):

explain it maybe? so i can understand for future questions /.\

hartnn (hartnn):

so you have to complete the square.

hartnn (hartnn):

To complete the square, add and subtract (co-efficient of 'x' /2)^2 first

hartnn (hartnn):

for x^2 - 8x + 13 = 0 what is the co-efficient of x ?

OpenStudy (anonymous):

1?

hartnn (hartnn):

for x^2 its 1 because x^2 =1x^2 what about 'x' ? which number is before 'x' (not x^2)

OpenStudy (anonymous):

oh 8

hartnn (hartnn):

close its -8 now next step is to divide it by 2

OpenStudy (anonymous):

divide what by 2

OpenStudy (anonymous):

sorry im slow /.\

hartnn (hartnn):

-8 "add and subtract (co-efficient of 'x' /2)^2" so add and subtract (-8 /2)^2 so add and subtract (-4)^2 = ... ?

OpenStudy (anonymous):

im confused :/

hartnn (hartnn):

with what ?

OpenStudy (anonymous):

-8-2 right?

OpenStudy (anonymous):

-8-2=-10?

hartnn (hartnn):

-8/2

OpenStudy (anonymous):

subtracting or dividing..............

hartnn (hartnn):

i never said subtracting by 2 its dividing by 2

OpenStudy (anonymous):

yea you did :/ up there but -4?

hartnn (hartnn):

yes, -4 now square it

OpenStudy (anonymous):

-4^2? is it the first answer choice

hartnn (hartnn):

whats the value of \(\large (-4)^2 =..\)

OpenStudy (anonymous):

what do you mean

hartnn (hartnn):

asking you the value of (-4)^2 (-4)*(-4) = ..

OpenStudy (anonymous):

16?

hartnn (hartnn):

yes so we add and subtract 16 \(x^2 -8x+16 + (13-16) = 0\)

hartnn (hartnn):

now \(x^2 -8x+16\) is the perfect square of \((x-4)^2 \) {because 16 is the square of 4}

hartnn (hartnn):

so we have \((x-4)^2 -3=0\)

hartnn (hartnn):

see the steps and ask doubts

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^